OpenStudy (anonymous):

3/7(y+2)=g solve for y

OpenStudy (zeesbrat3):

3/7y+6/7=g -6/7=-6/7 3/7y=g-6/7 x7/3=7/3 y=7/3g-2

OpenStudy (anonymous):

3/7y+6/7=g 3/7y=g-6/7 7/3 x 3/7y=g-6/7 x 7/3 y=g-2